OpenStudy (anonymous):

Express 40 = 1 using logarithmic notation. With explanation please.

OpenStudy (anonymous):

lol how is \(40=1\) ?

OpenStudy (anonymous):

here is my guess this really says \[4^0=1\] right?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Is it the same before? 4 = log[1] 0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[b^y=x\iff \log_b(x)=y\]

OpenStudy (anonymous):

log[4] (1)=0

OpenStudy (anonymous):

i mean it is the same as before, but your answer is wrong now it is right

OpenStudy (anonymous):

in fact \[\log_b(1)=0\] always, because \(b^0=1\) for any positive \(b\)

OpenStudy (anonymous):

Oh, Ok everything is all clear now. I can answer all the questions with no problem. Thank you so much.

OpenStudy (anonymous):

yw